OpenStudy (anonymous):

Solve 4/x^2-9=7/x^2-7x+12-5/x^2-x-12

OpenStudy (anonymous):

\[\frac{ 4 }{ x^2-9 }=\frac{ 7 }{ x^2-7x+12 }-\frac{ 5 }{ x^2-x-12 }\]

OpenStudy (anonymous):

what you could do is expand the terms on the bottom and then try to get common denominators and then you should be able to solve

OpenStudy (anonymous):

thank you, I got it and I was able to see the mistake I did. once again thank you for the help.